A) Predict whether the change in entropy (ΔS) is positive, negative or near zero for the following processes at 298K. Provide a molecular level explanation or justification for your prediction . (B) Now calculate the change in entropy (ΔS) for the following processes at 298K . Write a brief explanation on how your predicted vs. calculated results compare. i. H2O (s) → H2O (l), melting of ice. ii. Formation reaction for steam (water vapor, H2O (g)). iii. C3H8 (l) → C3H8 (aq), Liquid propane going to aqueous propane. iv. The hydrolysis of the dipeptide glycylglycine. All assumptions need to be clearly and concisely stated. If thermodynamic parameters are used, the citation, reference or link to where this thermodynamics data came from must be stated. The thermodynamic parameters are required for calculating entropy values. Prediction can be done with this data
Answer #1

Part A.

i) The change in entropy (ΔS) is positive.

Explanation: Here, ΔS = Sl - Ss, where Sl > Ss

Hence, ΔS is positive.

ii) The reaction for the formation of steam: H2(g) + 1/2 O2(g) ----> H2O(g)

The change in entropy (ΔS) is negative.

Explanation: Here, Δn=no. of gaseous moles of products - no. of gaseous moles of reactants = 1-(1 + 1/2) = -1/2

Hence, ΔS is also negative.

iii) The change in entropy (ΔS) is positive.

Explanation: Here, when liquid propane is converting to aqueous propane, the volume increases, as a result, the disorder/randomness does also increase.

Hence, ΔS is positive.

iv) The reaction: glycylglycine(s) + H2O(l) ----> Glycine(aq) + Glycine (aq)

The change in entropy (ΔS) is positive.

Explanation: glycylglycine is a solid dipeptide, which on hydrolysis produces two aqueous glycine species, which can make more disorder/randomness compared to the solid glycylglycine.

Hence, ΔS is positive.

Part B.

i) ΔS = SH2O(l) - SH2O(s) = 70 - 44.8 = 25.2 J/mol.K

ii) ΔS = SH2O(g) - (SH2(g) + 1/2 * SO2(g)) = 189 - (130.59 + 1/2 * 205.03) = -44.1 J/mol.K

iii) ΔS = SC3H8(aq) - SC3H8(l) = 340.2 - 171 = 169.2 J/mol.K

iv) ΔS = 2*Sglycine(aq) - (Sglycylglycine(s) =+ SH2O(l)) = 2*173.5 - (180.3 + 70) = 96.7 J/mol.K